# Investigating Resource-efficient Neutron/Gamma Classification ML Models Targeting eFPGAs

### v1

@@ 5 Conclusion
We have investigated the feasibility of deploying simple BDT and fcNN ML models on eFPGAs by using neutron/gamma classification as a case study and a stand-in commercial Artix 7-series FPGA target. We compared the study results to the golden standard of using the charge comparison method. The data set for the study was acquired with a fanout board where the choice of a 50 resistor in series with the SiPM dominates the single photon response shape. In this regime, we were able to show that an ADC with only 8-bits resolution and Nyquist sampling rates of approximately 15 MS/s can still allow us to maintain excellent neutron efficiency at very low gamma leakage values when using either multiple input feature ML architecture. We compare this against the standalone standard charge comparison metric. With the waveform digitizer option, a >= 10-bit, O(100) MS/s ADC has to be used to maintain comparable performance. With the analog integration option, a similar resolution ADC is required and sampling rate will be set by expected event rate requirements since only the final integration values are digitized.
More generally, we have shown that a low power, pipelined ML-aided solution is possible for real-time PSD functionality. An appropriately sized eFPGA fabric can be specified and synthesized to directly integrate on a single chip that includes direct SiPM readout front-ends and other required circuitry such as ADCs and TDCs. At 180nm, we expect the required area to be several mm by several mm to accommodate a properly sized eFPGA to implement the BDT models considered here. For fcNN models, the area required for most resource types will be smaller but additional area will be required for the necessary RAM blocks. If scaling down to 28nm, we expect the eFPGA fabric area requirements to decrease significantly.

### v2

@@ 5 Conclusion
[…]
We have calculated the resource types and quantities required for both ML architectures to implement non-trivial, high performance models to target the task of neutron/gamma classification. While the exact resource quantities and types differ, the general trade-off between BDT and fcNN models is that no DSP nor block RAM resources are required for BDT models but much larger numbers of LUTs are required compared to fcNN models. The translation to exact trade-offs in total required area will be process node and foundry dependent. However, having two fewer resource types to incorporate into the eFPGA fabric will simplify its design and implementation. The lower latencies and initiation intervals for BDTs provide an additional benefit if trying to maximize prediction throughput as is desired for the target applications for this case study. Thus, these two sets of benefits guide the decision to focus on BDT models and design the eFPGA fabric accordingly for a future first test chip. This first test chip will need to be larger than the one described in [7] if gamma leakage is required to be maintained. However, for an initial demonstration, relaxing this value to and optimizing the selection of input feature quantization can lower the amount of LUTs required to deploy high performance BDT models for neutron/gamma classification on eFPGA fabrics.